4. The table below presents macroeconomic data for two nations: Raljon and Kabukistan.

Raljon

Kabukistan

Nominal GDP growth

26%

6%

Inflation

20%

3%

Unemployment rate

4%

10%

Population growth

5%

1%

Assume that both nations start with real GDP of $1,000 per citizen. Fill in the blanks in the table below, assuming the data above applied for every year considered.

Raljon

Kabukistan

Real GDP growth rate per person

Years required for real GDP per person to double

Real per person GDP in 140 years
Answer #1

Raljon

Kabukistan

Real GDP growth rate per person

26-20-5 =1% 6-3-1 = 2%

Years required for real GDP per person to double

70/1 =70 70/2 = 35

Real per person GDP in 140 years